OpenStudy (he66666):

What is the derivative of y= 5^(-1/x)? The answer says y= 5^(-1/x) = [5^(-1/x)][In5][(-1)(-x^(-2)] = [5^(-1/x)(In5)]/x^2 but I don't get the second step "-x^(-2)".. why is x a negative, not positive? Since the exponent of the original question is (-1/x), shouldn't the inner derivative be [(-1)(x^(-2)] ?

OpenStudy (jamesj):

y = 5^(-1/x) implies that ln y = -1/x . ln 5 => y'/y = 1/x^2 . ln 5 => y' = ln 5 . 1/x^2 . y = ln 5 . 1/x^2 . 5^(-1/x)

OpenStudy (he66666):

Sorry, but can you put it in simpler terms (if possible)? We are learning about the chain rule and how to apply them...

OpenStudy (jamesj):

Write y =5^u where u = -1/x Then dy/dx = dy/du . du/dx Now du/dx = 1/x^2, yes? and dy/du = ln 5 . 5^u so far so good?

OpenStudy (he66666):

Yes, I'm following

OpenStudy (jamesj):

Then dy/dx = dy/du . du/dx = ln 5 . 5^u . (1/x^2) = ln 5 . 5^(-1/x). 1/x^2

OpenStudy (he66666):

Thank you! But I was wondering, if this is correct? u=-1/x = = -x^(-1). I can't seem to differentiate -x^(-1). I get dy/dx= (-1).(x^-2)

OpenStudy (jamesj):

If f(x) = 1/x = x^(-1), then df/dx = (-1)x^(-1-1) = -1/x^2 hence for our function u = -1/x, du/dx = -(-1/x^2) = 1/x^2

OpenStudy (he66666):

Oh okay. Thanks so much!